# Past, Present, and Future of Bug Tracking in the Generative AI Era

### v1

@@ 3.1.12. Patch Deployment
Patch deployment in contemporary software organizations is typically realized through CI/CD pipelines, supported by toolchains that automate the path from code to production. In practice, once a patch is verified and merged, CI/CD platforms such as Jenkins [^25] https://www.jenkins.io/, GitHub Actions [^26] https://github.com/features/actions, or GitLab [^27] https://about.gitlab.com/ CI trigger build processes, package the code, run integration and regression tests, and then promote the artifacts to staging or pre-production environments. From there, deployment into production may follow patterns such as blue–green deployments or canary releases, which reduce risk by gradually exposing the patch to subsets of users before a full rollout. Industry studies highlight that these practices significantly accelerate release frequency and improve reliability. For instance, case evidence from Paddy Power’s adoption of continuous delivery demonstrates that releases shifted from a few times per year to weekly or even daily, with deployment becoming a single-button operation instead of a multi-day manual task, leading to dramatic improvements in productivity, quality, and customer satisfaction (Chen, 2015). Likewise, a multiple-case study of Finnish software-intensive companies shows that organizations with more automated toolchains, especially those minimizing manual steps, are able to deploy faster and more reliably, while gaps in automation—such as missing performance testing or acceptance testing—directly slow down the delivery cycle (Mäkinen et al., 2016).
@@ 3.2. Architecture of the Proposed Bug Tracking System
Our proposed bug tracking system employs a multi-LLM agent design in which specialized LLM agents are assigned to discrete steps in bug tracking. Rather than relying on a single monolithic model to handle all tasks, each agent is optimized and prompted for a particular subtask and consumes a curated context tailored to that role, such as historical bug fixes for localization and no-code fix suggestions. This decomposition enables each agent to use prompts and toolchains best suited to its objective, retain compact, task-relevant context for efficiency, and produce more focused, explainable outputs that can be validated in an HIL setting. Agents communicate via a lightweight coordination layer so that the outputs of upstream agents can be used by downstream agents. For example, the enhanced bug report becomes the canonical input for downstream agents like bug localization and patch-generation agents. Throughout the workflow, HIL supervisors (customer support staff, test engineers, PMs, or developers) oversee critical handoffs and validations.
While managing this complex system, which contains multiple data sources, toolkits, and multiple agents, we adopt the Model Context Protocol (MCP) (Krishnan, 2025) pattern. A dedicated MCP server acts as a broker and context store, exposing interfaces for prompt templates, cached retrieval vectors for RAG, historical artifacts such as old bug reports, project docs and wiki, developer lists, and source code, and tool integrations like the reproduction toolkit and ML models. The MCP centralizes agent registration, versioning, and policy enforcement while enabling agents to remain lightweight and interchangeable.
Figure 3 visualizes this architecture. On the left, persistent knowledge sources and tooling are shown (from the client’s initial bug report to prompt libraries). The MCP server in the middle mediates access to these resources and hosts the coordination logic. To the right, there is a sequence of LLM agents; each row corresponds to one agent and its primary artifact, for example, LLM Agent for Bug Report Creation and Bug Report. Arrows indicate data flow: the initial report and auxiliary resources are read by the creation agent, the enhanced report is produced and sent to the MCP server and stored for downstream agents, and intermediate outputs like classification and validation results are stored for consumption by later stages. At the bottom, a deployment automation tool receives a verification result and produces a deployment report, closing the pipeline. We removed the MCP clients that exist in each MCP connection for the sake of simplicity from Figure 3. The diagram intentionally separates agents from resources and shows MCP as the single coordination and integration point, highlighting the auditability and reproducibility of every step.
In summary, the architecture provides a structured way to operationalize multiple LLM agents within a bug tracking system. By combining agent specialization with a central MCP for orchestration, we gain modularity, reproducibility, and governance. Agents can be updated, rolled back, or replaced independently; provenance for every recommendation is recorded; and HIL checkpoints are enforced where human judgment is required. Remaining system-level challenges include ensuring consistency of shared state across agents, managing prompt and model drift, verifying agent outputs at scale, and enforcing security and privacy policies. Addressing these will be key in future work to make multi-LLM bug tracking systems robust and production-ready.
@@ 3.3. Roles and Responsibilities of Each Stakeholder
Table 2. Transformation of Roles and Responsibilities of Stakeholders From Current Bug Tracking Workflow to Proposed Workflow
@@
Ops Team Manually deploys patches and monitors the production environment. Focuses on infrastructure and CI/CD pipelines due to automated deployment and testing.
@@
Developer: In the current settings, developers manually localize and reproduce assigned bugs. After that, they implement code fixes. In the proposed workflow, developers act as supervisors in the HIL patch generation step. LLM agents handle bug reproduction and localization, and suggest code patches. Developers review, refine, and approve these changes, ensuring correctness and maintainability. They remain ultimately responsible for code quality.
@@
Tester: In the current workflow, testers manually create and execute test cases to verify code fixes. In the proposed workflow, testers act as supervisors in the HIL testing step. Test agents generate test suites, execute them, and verify both code fixes. Testers oversee this process, review generated test cases, and add or refine tests that exceed the LLMs’ current capabilities. Their role shifts from execution to validation and augmentation, effectively becoming “Test Reviewers.”
Ops Team: In the current bug tracking setup, the operations team manually deploys patches after verification and monitors the production environment. In the proposed workflow, the ops team’s role becomes supervising LLM agents that take care of the ops team’s tasks. LLM agents develop CI/CD pipelines and maintain them, while the ops team supervises and makes necessary modifications.
@@ 4. Discussion
This section critically discusses the implications of the proposed system, exploring its potential impact and the directions for future research in this field. Section 4.1 outlines major challenges and risks that must be addressed, ranging from technical hurdles like cascading errors and limited generalization to broader concerns regarding accountability and economic costs. In Section 4.2, the discussion turns to the practical implications for practitioners, detailing how the system’s modularity, HIL design, and toolchain integration can be leveraged to transform software development workflows. Section 4.3 concludes the section by examining the implications for researchers, identifying a new landscape of research questions concerning task ordering, human oversight, and the evaluation of complex agent-based systems.
@@ 4.1. Future Challenges and Risks
4.1.1. Accumulated Errors Due to Multi-Step LLM Dependency
One of the primary risks in our approach is that multiple stages of the bug lifecycle, such as reporting, reproduction, localization, and patch generation, depend on LLMs, each of which has inherent limitations in accuracy. Since each step builds upon the output of the previous step, poor performance may propagate and amplify, similar to earlier studies (Jimenez et al., 2023). This cascading effect could significantly reduce the final accuracy of bug resolution.
@@ 4.1.2. Accountability Issue
4.1.3. Bias and Inaccuracy in LLM Predictions
LLMs and ML models are trained on large-scale datasets that can contain biases or knowledge gaps, leading to misclassification of bugs, misleading explanations, or incorrect fixes caused by various factors such as code generation mistakes (Chen et al., 2024). In specialized software domains, where data is scarce or domain knowledge is highly specific, LLMs may struggle to provide accurate results. The challenge lies in determining the extent of these biases and inaccuracies and understanding how they affect different stages of bug tracking and resolution.
@@ 4.1.4. Limited Generalization Across Different Software Projects
A major future challenge in AI-driven bug tracking is ensuring that models generalize well across diverse software ecosystems. The effectiveness of an LLM-based bug tracking system may vary significantly depending on the programming languages, frameworks, and architectures involved. A model trained on one software type may struggle when applied to another, leading to inconsistencies in bug resolution accuracy. Future research must explore how to improve adaptability across different software contexts and what factors influence the performance of LLMs in varied environments.
4.1.5. Evaluation

### v2

@@ 3.1.12. Patch Deployment
Patch deployment in contemporary software organizations is typically realized through CI/CD pipelines, supported by toolchains that automate the path from code to production. In practice, once a patch is verified and merged, CI/CD platforms such as Jenkins [^25] https://www.jenkins.io/, GitHub Actions [^26] https://github.com/features/actions, or GitLab [^27] https://about.gitlab.com/ CI trigger build processes, package the code, run integration and regression tests, and then promote the artifacts to staging or pre-production environments. From there, deployment into production may follow patterns such as blue–green deployments or canary releases, which reduce risk by gradually exposing the patch to subsets of users before a full rollout. Industry studies highlight that these practices significantly accelerate release frequency and improve reliability. For instance, case evidence from Paddy Power’s adoption of continuous delivery demonstrates that releases shifted from a few times per year to weekly or even daily, with deployment becoming a single-button operation instead of a multi-day manual task, leading to dramatic improvements in productivity, quality, and customer satisfaction (Chen, 2015). Likewise, a multiple-case study of Finnish software-intensive companies shows that organizations with more automated toolchains, especially those minimizing manual steps, are able to deploy faster and more reliably, while gaps in automation, such as missing performance testing or acceptance testing, directly slow down the delivery cycle (Mäkinen et al., 2016).
@@ 3.2. Architecture of the Proposed Bug Tracking System
Our proposed bug tracking system employs a multi-LLM agent design within a structured layered architecture. In this design, specialized LLM agents are assigned to discrete steps in the bug tracking lifecycle. Rather than relying on a single monolithic model to handle all tasks, each agent is optimized and prompted for a particular subtask, such as historical bug fixes for localization or no-code fix suggestions. This decomposition enables each agent to use toolchains best suited to its objective, retain a compact, task-relevant context for efficiency, and produce more focused, explainable outputs. To ensure seamless interoperability between these specialized components, agents communicate via a standardized Agent-to-Agent (A2A) protocol (Surapaneni et al., 2025), which defines the contract for data exchange and state handoffs.
To manage the complexity of multiple data sources, toolkits, and agents, we adopt the Model Context Protocol (MCP) (Krishnan, 2025) pattern. This approach centralizes context management and tool integration, allowing the agent microservices to remain lightweight and interchangeable while ensuring consistent access to external resources and LLM APIs.
Figure 3 visualizes this layered architecture, organized into five distinct horizontal levels:
  1. (1)
    External Systems Layer: Located at the top, this layer handles interactions with the outside world, including the Bug Reporting Client for user inputs and the LLM API Provider that powers the intelligence of the system.
  2. (2)
    Platform Orchestration Layer: This layer serves as the system’s central nervous system. It contains the Workflow Orchestrator, which manages the sequence of agent execution and state transitions adhering to the A2A protocol. It orchestrates the flow of control and data between the upper client layers and the specific agent microservices.
  3. (3)
    Agent Microservices Layer: This layer hosts the sequence of specialized agents operating under the A2A protocol. The workflow begins with the Bug Report Creation and Enhancement Agents, moving through Reproduction, Classification, and Validity Check Agents. Subsequently, the pipeline triggers the resolution phase with the No-Code Fix, Bug Localization, Developer Assigner, Patch Generation, and Bug Fix Verification Agents. Successful verification leads to the external Deployment Automation stage.
  4. (4)
    MCP Layer: Situated centrally between the agents and the data stores, this layer functions as the integration broker. The MCP Server mediates access between the agents in the layer above and the external systems or data layers below. It standardizes prompt templates and tool integrations, while managing the connection to the LLM API Provider to supply the agents with inference capabilities.
  5. (5)
    Data Persistence Layer: At the bottom, the architecture utilizes specialized storage solutions to maintain system state and knowledge. A Relational DB stores structured data such as bug reports, developer lists, and metadata. A Vector DB handles semantic data, including code embeddings, project documentation, and wikis. Finally, Blob Storage is used for unstructured artifacts like screenshots, patch files, and verification results.
In summary, the architecture provides a structured way to operationalize multiple LLM agents. By organizing the system into logical layers—separating orchestration, agent logic, context management, and data persistence—we gain modularity, scalability, and governance. The A2A protocol, combined with the distinct MCP layer, ensures that, despite the distributed nature of the microservices, data access, inter-agent communication, and tool usage remain consistent and auditable. Remaining system-level challenges include ensuring state consistency across the orchestration layer and managing the latency of sequential agent interactions. Addressing these will be key in future work to make multi-LLM and multi-agent-powered bug tracking systems robust and production-ready.
3.3. Early Adaptations of Proposed Framework
To provide feasibility-oriented empirical evidence for our proposed framework, we report results from a few implemented subcomponents, including modules introduced and evaluated in our own prior work. We first summarize empirical findings on early stages that are central to real-world bug tracking, bug report enhancement, reproduction from natural-language reports, and validity detection. We then situate these findings alongside representative recent systems that empirically demonstrate the feasibility of downstream repair stages such as localization, patch generation, and verification.
In the domain of bug report enhancement, ”ImproBR” is introduced, which is an agentic framework designed to transform ambiguous, user-generated bug reports into structured, actionable documents by refining critical sections such as S2R, OB, and EB  (Emre Furkan Akyol and Tuzun, 2026). The methodology employs a multi-stage pipeline that first utilizes a fine-tuned DistilBERT classifier and heuristic analysis to detect missing or low-quality sections, followed by a ”Bug Report Improver Agent” powered by GPT-4o mini. To ensure domain accuracy and minimize hallucinations, the agent leverages RAG to access knowledge from the Minecraft Wiki.The evaluation of ImproBR was conducted on the Mojira datasetResults demonstrate substantial improvements in both structure and content quality. First, the system increased the average structural completeness of reports from 7.9% to 96.4% by automatically generating missing S2R, OB, EB sections. In terms of content richness, the average word count for S2R sections rose from 9.1 to 111.9, and the usage of domain-specific terminology increased from 0.36 to 4.59 per report. Furthermore, the semantic similarity of generated S2R sections to ground-truth duplicate reports improved from 0.020 to 0.180, indicating a closer alignment with expert-written content. Crucially, a manual evaluation of 139 challenging, real-world bug reports confirmed the practical impact of these enhancements. ImproBR more than doubled the proportion of executable S2R sections, increasing the average from 28.8% to 67.6%, and successfully raised the number of fully reproducible bug reports from 1 to 13.
Focusing on bug reproduction, ”BugCraft” is introduced, an end-to-end automated framework capable of reproducing crash bugs in the complex, open-ended environment of Minecraft directly from user descriptions (Yapağcı et al., 2025). The system’s methodology comprises two distinct stages: a “Step Synthesizer” that uses LLMs augmented with RAG to convert unstructured reports into high-quality, clustered S2R, and an ”Action Model” vision-based agent leveraging a custom macro API and OmniParser to execute these steps within the game client. The system is evaluated on ”BugCraft-Bench,” a curated dataset of 86 reproducible crash reports, where BugCraft achieved a state-of-the-art success rate of 34.9% using GPT-4.1, significantly outperforming general computer-use baselines like OpenAI’s CUA (25.5%) and UI-TARS (0%). Additionally, the system proved highly efficient, reducing the cost per attempt to approximately \$1.16 compared to the significantly higher costs of manual reproduction, as well as a decrease in TTR from manual 3.41 days to automated 10 minutes.
Regarding the bug report validity detection, we conducted a comprehensive comparative study titled ”Judge the Votes” to assess the efficacy of various models in classifying bug reports as valid or invalid (Dinç and Tüzün, 2025). The methodology involved training and evaluating a spectrum of models including classical ML algorithms (SVM, Random Forest), fine-tuned transformers (RoBERTa, BERT), and reasoning LLMs (GPT-o3-mini, Llama-3) using a dataset of 10,000 Firefox bug reports; we also introduced a novel ”Judge LLM” ensemble that combines the probabilistic votes of classical classifiers with RAG-retrieved few-shot examples to render a final decision and explanation. Our results established that fine-tuned semantic classifiers are superior, with RoBERTa achieving the highest $F_{1}$ score of 0.909, whereas standalone LLMs lagged behind but showed dramatic improvement when augmented with RAG (rising to an $F_{1}$ of 0.815). Uniquely, the Judge LLM pipeline achieved a competitive $F_{1}$ of 0.871 while providing valuable natural language explanations for its verdicts.
In addition to our own early implementations, several recent LLM-based automated software engineering frameworks provide strong empirical evidence that the core components of our proposed framework are already technically feasible at scale. While some of these systems adopt fully agentic control loops and others deliberately avoid persistent agent state, together they demonstrate that complex multi-stage bug resolution pipelines can already be constructed with current LLM-based technologies.
Agentless (Xia et al., 2025) presents a repository-level automated bug-fixing system that intentionally avoids maintaining persistent environment state or long-horizon agent memory. Instead, it relies on structured reasoning, static repository analysis, and targeted external tool calls. The system decomposes bug fixing into four major stages: issue understanding, file-level localization, patch generation, and verification. For localization, Agentless applies hierarchical reasoning over repository structure and file contents; for patch generation, it uses LLM-guided program transformation; and for validation, it executes candidate patches against available test suites. On the SWE-bench benchmark, Agentless achieves a verified fix rate of 27.8%, substantially outperforming many prior systems that rely on heavier infrastructure. Notably, the system maintains high efficiency, solving tasks with significantly fewer tool calls and reduced runtime overhead. These results demonstrate that even without a fully agentic architecture, sophisticated multi-stage bug resolution pipelines are already realizable, directly supporting the feasibility of our framework’s design.
AutoCodeRover (Zhang et al., 2024) proposes a fully automated program repair agent built around an iterative generate–execute–repair loop. The system first performs hierarchical localization, identifying suspicious files and functions using repository analysis and LLM-guided reasoning. It then synthesizes candidate patches and validates them through execution-based testing, refining unsuccessful patches across multiple iterations. On the SWE-bench-Verified benchmark, AutoCodeRover achieves a state-of-the-art verified fix rate of 46.2%. In ablation studies, removing execution-based validation drops performance by over 15 percentage points, highlighting the importance of tight integration between generation and verification. AutoCodeRover also demonstrates strong efficiency: for many tasks, it converges to correct patches within 3–6 iterations, showing that agentic feedback loops can quickly stabilize toward valid solutions. These findings closely mirror the patch generation and verification stages of our proposed framework.
OpenHands (Wang et al., 2024e) presents a general-purpose software engineering agent designed to execute complex multi-step development workflows. Its architecture integrates planning, tool usage, repository navigation, and execution feedback into a unified decision-making loop. The CodeActAgent variant performs best, combining code generation with explicit action planning and execution. On USEbench, a benchmark consisting of 1,271 real-world repository-level tasks, CodeActAgent achieves an overall task success rate of 26.8%. On software maintenance subsets of the benchmark, including bug fixing and feature implementation, success rates exceed 30%. The system also demonstrates strong generalization across programming languages and project scales, providing concrete evidence that LLM-driven agents can coordinate complex end-to-end software workflows similar to those envisioned in our framework.
USEagent (Applis et al., 2025) proposes a unified agent architecture capable of handling a broad spectrum of software engineering tasks including bug fixing, regression testing, code generation, and test generation. The agent integrates hierarchical planning, long-horizon reasoning, repository exploration, and execution feedback into a single control loop. On USEbench, USEagent achieves 33.3% overall task efficacy, significantly outperforming OpenHands (26.8%). On software maintenance tasks derived from SWE-bench-Verified, USEagent reaches 45.6% verified fix rate, approaching the specialized performance of AutoCodeRover (46.2%), while maintaining substantially broader task coverage. Importantly, USEagent exhibits strong stability: over 70% of successful tasks converge within five agent iterations, indicating that agentic workflows can remain both scalable and predictable in practice.
While the above systems demonstrate impressive capabilities in automated localization, patch generation, and verification, they all operate under the assumption that the bug is already well-formed, reproducible, and ready for repair. In contrast, our framework is explicitly designed to handle the practical challenges of real-world bug tracking, where bug reports are often incomplete (Zhang et al., 2017), ambiguous (Bettenburg et al., 2008), invalid (Wu et al., 2020), or non-reproducible (Goyal and Sardana, 2019). Rather than assuming perfect inputs, our system actively improves bug reports through interactive enhancement, attempts automated reproduction from natural language descriptions, classifies bugs, performs bug–feature tracing, supports no-code fixes for invalid reports, assigns bugs with accountability constraints, and integrates patch deployment and user verification into a unified lifecycle. By covering the entire pipeline from raw user submission to final resolution and feedback, our framework addresses substantial portions of the bug lifecycle that existing systems intentionally leave outside their scope.
Taken together, these systems demonstrate that the major components of our proposed bug tracking framework, automated localization, patch generation, iterative verification, multi-stage reasoning, and workflow orchestration, are already realizable with current agentic technologies. When combined with our own early results on bug report enhancement, bug reproduction, and validity detection, these findings provide strong empirical grounding for the feasibility of an end-to-end AI-powered bug tracking framework.
In the commercial landscape, the shift from passive issue tracking to AI-native automated resolution is already underway. A new generation of AI-powered debugging and maintenance platforms provides industrial validation for different phases of our proposed framework. Here, we show some of the recent developments in the industrial landscape.
Regarding the creation and enhancement of bug reports, industry tools have largely deprecated the manual entry of environment details. Platforms such as BetterBugs [^28] https://www.betterbugs.io, Jam [^29] https://jam.dev, Bugpilot [^30] https://www.bugpilot.com/, and LogRocket [^31] https://logrocket.com/ have operationalized the concept of a bug creation agent. These tools automatically capture comprehensive telemetry, including console logs, network requests, device specifications, and session replays, at the moment of failure, eliminating the need for back-and-forth communication between users and developers. For instance, Bugpilot combines user-side reporting with automated error monitoring to capture ”hidden bugs” that users fail to report, while Jam integrates directly with issue trackers like Jira to inject this rich context into the developer’s existing workflow.
For bug reproduction and localization, the industry is moving toward deterministic replay and automated test generation. Replay.io  [^32] https://replay.io/ and rrweb [^33] https://www.rrweb.io/ record execution traces (DOM updates, network events) to render bugs deterministically replayable. Similarly, QA Wolf  [^34] https://www.qawolf.com/ demonstrates the viability of transforming user flows into structured, maintainable end-to-end test suites, effectively bridging the gap between a bug report and a reproducible test case.
Finally, regarding bug fixing, several platforms are deploying agents that mirror the Bug Fix and Verification components of our framework. TraceRoot.AI [^35] https://traceroot.ai/ and Keystone [^36] https://www.withkeystone.com/ connect directly to observability data (logs and traces) to construct execution trees, identify root causes, and automatically generate draft PRa to resolve production issues. Jazzberry.ai [^37] https://www.jazzberry.ai/ further validates the verification aspect by executing code changes in secure, sandboxed microVMs to detect runtime errors before they reach production.
While these commercial solutions demonstrate the industrial viability of individual agentic capabilities, they currently lack rigorous results to be presented here, and they exist as fragmented point solutions, each handling either reporting, reproduction, or fixing in isolation. Our proposed framework unifies these disparate capabilities into a single, cohesive multi-agent orchestration, managing the entire lifecycle from the initial user report to the verified deployment of the fix.
@@ 3.4. Roles and Responsibilities of Each Stakeholder
Table 2. Transformation of Roles and Responsibilities of Stakeholders From Traditional Bug Tracking Workflow to Proposed Workflow
@@
Ops Team Oversees automated deployment and CI/CD pipelines. Oversees automated deployment with the assistance of LLM agents.
@@
Developer: In the traditional settings, developers manually localize and reproduce assigned bugs. After that, they implement code fixes. In the proposed workflow, developers act as supervisors in the HIL patch generation step. LLM agents handle bug reproduction and localization, and suggest code patches. Developers review, refine, and approve these changes, ensuring correctness and maintainability. They remain ultimately responsible for code quality.
@@
Tester: In the traditional workflow, testers manually create and execute test cases to verify code fixes. In the proposed workflow, testers act as supervisors in the HIL testing step. Test agents generate test suites, execute them, and verify both code fixes. Testers oversee this process, review generated test cases, and add or refine tests that exceed the LLMs’ current capabilities. Their role shifts from execution to validation and augmentation, effectively becoming “Test Reviewers.”
Ops Team: In the traditional setup, the operations team manages automated deployment and oversees CI/CD pipelines. In the proposed workflow, their core responsibilities remain largely unchanged as deployment is already automated. The primary distinction is that the ops team now oversees the deployment process with the assistance of LLM agents, utilizing them for monitoring and anomaly detection while maintaining final authority over the pipeline.
@@ 4. Discussion
This section provides a critical synthesis of the proposed agentic framework, exploring its capacity to redefine bug tracking while identifying the multifaceted barriers to its realization. Section 4.1 examines a comprehensive range of future challenges and risks, encompassing technical failure modes like hallucination and cascading errors, ethical and regulatory hurdles such as bias and privacy, and the inherent difficulty of managing non-deterministic bugs. Section 4.2 translates these findings into practical implications for practitioners, detailing how modularity and HIL design can be leveraged to optimize industrial workflows. Finally, Section 4.3 charts a course for future research, identifying a new landscape of inquiries concerning pipeline orchestration and human-agent accountability.
@@ 4.1.2. Limited Generalization Across Different Software Projects
A major future challenge in AI-driven bug tracking is ensuring that models generalize well across diverse software ecosystems. The effectiveness of an LLM-based bug tracking system may vary significantly depending on the programming languages, frameworks, and architectures involved. A model trained on one software type may struggle when applied to another, leading to inconsistencies in bug resolution accuracy. To alleviate this limitation, future systems can support cross-project adaptation by dynamically constructing task-specific model context from relevant project artifacts, such as source code, configuration files, documentation, historical bug reports, and execution traces. Instead of relying on static model knowledge or retraining, the framework can assemble only the most relevant project information at runtime and inject it into the model’s working context. This allows the system to adapt to new languages, frameworks, and architectures while preserving generalization across heterogeneous software ecosystems.
4.1.3. Bias in LLM Predictions
LLMs and ML models are trained on large-scale datasets that inevitably encode social, cultural, and technical biases as well as systematic knowledge gaps. These biases, when LLMs are used in bug tracking workflows, may propagate into downstream activities such as severity assessment and report classification, potentially causing uneven treatment of bugs across platforms or programming languages (Chen et al., 2024). In specialized software domains, where training data is scarce or domain knowledge is highly specific, such biases may become more pronounced. The challenge lies in identifying how these embedded biases influence different stages of bug tracking and understanding their impact on the fairness and reliability of downstream decisions. Such limitations reflect fundamental constraints of model training and data representation, affecting the baseline trustworthiness of individual predictions produced by the system. To mitigate the impact of such biases in future deployments of our framework, continuous bias auditing and calibration mechanisms can be employed to monitor skewed predictions over time. Moreover, placing human oversight at critical decision points, such as bug validation and assignment, provides an additional safeguard against biased or systematically skewed predictions before they propagate through the pipeline.
4.1.4. Factuality Challenges and Hallucination
A major reliability challenge in LLM-powered bug tracking systems arises from two closely related but conceptually distinct failure modes: factuality errors and hallucination. Although these phenomena are often conflated in practice, recent work explicitly distinguishes between them. In particular, Bang et al. (Bang et al., 2025) characterize hallucination as model outputs that are inconsistent with either the model’s pretraining corpus or the provided input context, while factuality concerns the correctness of generated content with respect to an external ground-truth oracle. As a result, a response may be factually incorrect without being hallucinated if it remains consistent with the model’s internal knowledge, and conversely, a response may be hallucinated even when it happens to be factually correct.
Within automated bug tracking workflows, this distinction becomes operationally important. Factuality challenges arise when agents generate explanations, reproduction steps, environment assumptions, or patch logic that conflict with the actual behavior of the target system, underlying libraries, or current software ecosystem (Alansari and Luqman, 2025). These errors are particularly problematic in rapidly evolving software domains, where the real-world knowledge landscape changes after model training, causing internal model knowledge to become outdated and increasing the risk of factual failures (Huang et al., 2025). Hallucination, by contrast, manifests as the generation of plausible but unsupported code artifacts and behaviors, including fabricated functions, incorrect APIs, and semantically invalid program logic (Zhang et al., 2025a). In a bug tracking context, this may appear as nonexistent configuration settings, invalid execution steps, or misleading code modifications that conflict with the bug report or repository state. Both failure modes pose serious risks in high-stakes tasks such as bug report enhancement, reproduction, and patch generation, where confidently incorrect outputs can mislead downstream automation and human reviewers.
The severity of these risks is reinforced by recent large-scale analyses. Multiple surveys demonstrate that hallucination is not merely a surface-level generation artifact but an inherent limitation of current LLM generation mechanisms, emerging from probabilistic next-token prediction under uncertainty and imperfect contextual grounding (Xu et al., 2024; Huang et al., 2025). These studies further emphasize that both hallucination and factuality errors systematically undermine trustworthiness in real-world applications requiring correctness and verifiability (Alansari and Luqman, 2025; Huang et al., 2024b). Addressing both failure modes is therefore essential for reliable agentic bug tracking.
Prior works propose a range of mitigation strategies. RAG improves factual grounding by anchoring model outputs in external evidence, substantially reducing both factual errors and unsupported fabrications (Huang et al., 2025, 2024b). More recently, Self-Reflective Retrieval-Augmented Generation (SELF-RAG) extends this idea by allowing the model to dynamically decide when retrieval is needed and to critique both retrieved evidence and its own generations through explicit self-reflection tokens, leading to significant improvements in factuality and citation accuracy across diverse tasks (Asai et al., 2024). Self-reflection and iterative reasoning mechanisms further improve reliability by enabling models to detect and revise their own errors across multiple passes (Ji et al., 2023). Additional techniques, including hallucination-aware fine-tuning, calibration, and auxiliary verification modules, have also demonstrated effectiveness in reducing hallucinated content, particularly in code generation settings (Zhang et al., 2025b).
In our framework, these model-level techniques can be complemented by system-level safeguards. Each major pipeline stage could be subject to validation and human oversight, allowing low-confidence or low-quality agent outputs to be detected early. When such behavior is observed, the pipeline may be halted and responsibility would be transferred to human operators, preventing erroneous outputs from propagating across subsequent stages and preserving the integrity of the overall bug resolution process. This layered mitigation strategy would enable both factuality challenges and hallucination to be addressed through coordinated model improvements and robust system-level governance.
4.1.5. Accumulated Errors Due to Multi-Step LLM Dependency
One of the primary risks in our approach is that multiple stages of the bug lifecycle, such as reporting, reproduction, localization, and patch generation, depend on LLMs, each of which has inherent limitations in accuracy. Since each step builds upon the output of the previous step, poor performance may propagate and amplify, similar to earlier studies (Jimenez et al., 2023). Thus, even small inaccuracies introduced early in the pipeline could compound into major failures at later stages, significantly reducing the final accuracy of bug resolution.
To address this risk, our framework could employ explicit error detection and propagation control mechanisms throughout the pipeline. Each major stage could produce, in addition to its primary output, an associated confidence signal derived from model uncertainty, execution feedback, or downstream validation results.
For example, during bug reproduction, an agent could attempt to execute the synthesized S2R in a sandboxed environment; repeated execution failures, unexpected exceptions, or inconsistent outputs across multiple runs could be interpreted as strong indicators of low confidence in the reproduction result. Similarly, during patch generation and verification, failing test cases, compiler errors, or unstable behavior under repeated execution could automatically lower confidence in the generated patch.
When such confidence falls below predefined thresholds, automated processing could be halted, and the task could be escalated to human operators for inspection and correction. Furthermore, validation checkpoints between stages could prevent unreliable intermediate artifacts, such as incorrect reproduction traces or unstable patches, from propagating further across the workflow. By combining confidence-aware gating, human-in-the-loop escalation, and stage-wise validation, the framework could limit cascading failures while preserving the efficiency benefits of automation.
4.1.6. Agentic Evaluation Challenges